x/9 -y/7= -1
x-y=-3
solve the system with the addition method

Answer by checkley77(12844)

x/9-y/7=-1 MULTIPLY BY -7 & ADD.
x-y=-3
-7X/9+Y=7
X-Y=-3
-------------------
X-7X/9=4
(9X-7X)/9=4
2X=36
X=36/2
X=18 ANS.
18-Y=-3
-Y=-3-18
-Y=-21
Y=21 ANS.